The fundamental frequency on a aluminum string that is fixed at both ends is 468 Hz. The string is then cooled 174°C and as such its length changes. Determine how much the fundamental frequency will change as a result assuming the tension remains constant.
Δf1 =
1
Expert's answer
2020-04-20T10:41:30-0400

The fundamental frequency of string that is fixed at both ends

"f_1=\frac{v}{2L}."

Since the tension remains constant, the speed of wave remains constant too.

So

"f_1'=\frac{v}{2L'}."

The linear thermal expansion of string is described by equation

"L'=L(1+\alpha \Delta T), \; \alpha=25\times 10^{-6}\frac{1}{\degree \rm C}"

Hence, the change of fundamental frequency due thermal expansion of string

"\Delta f_1=\frac{v}{2L(1+\alpha \Delta T)}-\frac{v}{2L}=-f_1\alpha \Delta T"

"\Delta f_1=-468\times 25\times 10^{-6}\times (-174)=2.0\:\rm Hz."